OpenStudy (anonymous):

Five students took and algebra test. The average score was 84. The four scores were 96, 68, 79, and 85. What score did the fifth student get?

OpenStudy (anonymous):

(96+68+79+85+x)/5 = 84 (96+68+79+85+x)= 84*5 328+x =420 so, x= 92

OpenStudy (anonymous):

so fifth score is 92

OpenStudy (anonymous):

remember that the average is the sum of the terms divided by the number of the times. Since the number of terms is 5. we have 84=total/5 thus total=420 and 420= 96+68+79+85+x = 328+x and so x=92

OpenStudy (anonymous):

thank you everyone. :)